OpenStudy (anonymous):

Rationalize the denominator and simplify. (√(a+1)-2)/(√(a+1)+2)

OpenStudy (radar):

\[\sqrt{a+1}-2\over \sqrt{a+1}+2\]Multiply numerator and denominator by\[\sqrt{a+1}-2\]getting:\[a-4\sqrt{a+1}+5\over a-3\]